Given that VS = 3V, VBE = 0.3V, IC = 9mA, and R1 = 1kΩ, calculate the base current of the transistor Q1 using KVL. Also find the current gain (beta ) of the transistor Q1.


1
Expert's answer
2021-08-22T15:47:11-0400

We know that load line equation

"V_{BB}=V_{BE}+I_BR_B\\V_{BB}=V_S\\V_{S}=V_{BE}+I_BR_B"

Put value

"3=0.3+I_B\times1000"

"2.7=I_B\times1000"

"I_B=2.7mA"

"I_c=9mA"

"\beta=\frac{I_c}{I_B}=\frac{9\times10^{-3}}{2.7\times10^{-3}}=" 3.33